OpenStudy (anonymous):

Can someone people help me on this:?! Evaluate: I will put them in comments!

OpenStudy (anonymous):

1. \[\sqrt[3]{-27}\] 2. \[\sqrt{(-5a)^2}\]

OpenStudy (anonymous):

first one, look for the number whose cube is \(-27\), i.e. think of an \(x\) where \(x^3=-27\) the answer should come to you

OpenStudy (anonymous):

second one, square root of anything squared is the absolute value of anything , i.e. \[\sqrt{x^2}=|x|\]